# Multimodal Data Integration for Oncology in the Era of Deep Neural Networks: A Review

### v2

@@ 5 Challenges in Multimodal Learning (MML)
5.1 Availability of Large Amounts of High-quality Data
DL models are traditionally trained on large datasets with enough samples for training, validation, and testing, such as JFT-300M[147] and YFCC100M[148], which are not available in the cancer domain. For example, the largest genomics data repository, Gene Expression Omnibus (GEO) database, has approximately 1.1 million samples with the keyword ‘cancer’ compared to 3 billion images in JFT-300M [149]. Annotating medical and oncology data is a time-consuming and manual process that requires significant expertise in many different areas of medical sciences. Factors like heterogeneity of the disease, noise in data recording, background and training of medical professionals leading to inter- and intra-operator variability, cause lack of reproducibility and inconsistent clinical outcomes [150].
@@ 5.2 Data Registration and Alignment
Data alignment and registration refer to the process of combining and aligning data from different modalities in a useful manner [151]. In multimodal oncology data, this process involves aligning data from multiple modalities such as CT, MRI, PET, WSIs images, as well as genomics, transcriptomics, and clinical records. Data registration involves aligning the data modalities to a common reference frame and may involve identifying common landmarks or fiducial markers. If the data is not registered or aligned correctly, it may be difficult to fuse the information from different modalities accurately [152].
@@ 5.3 Pan-Cancer Generalization and Transference
Transference in MML aims to transfer knowledge between modalities and their representations to improve the performance of a model trained on a primary modality [152]. Because of the unique characteristics of each cancer type and site, it is challenging to develop models that are able to generalize across different cancer sites. Furthermore, the models trained on a specific modality, such as radiology images, will not perform well with other imaging modalities, such as histopathology slides. Fine-tuning the model on a secondary modality, multimodal co-learning, and model induction are a few techniques to achieve transference and generalization [153]. To overcome this challenge, there is a need to devise mechanisms for improved universality of ML models.
@@ 5.4 Missing Data Samples and Modalities
The unavailability of one or more modalities or absence of samples in a modality affects the model learning, as most of the existing DL models cannot process the “missing information”. This requirement, in turn, constrains the already insufficient size of datasets in oncology. Almost all publicly available oncology datasets have missing data for a large number of samples [149]. Various approaches for handling missing data samples and modalities are gradually gaining traction, however, this is still an open challenge in oncology [154].
@@ 5.5 Imbalanced Data
Class imbalance refers to the situation where one class (e.g., cancer negative/positive) is significantly more frequent in the data than another class. Class imbalance is a common issue in oncology data [154]. DL models struggle to accurately classify underrepresented classes. Techniques such as data augmentation, ensemble, continual learning, and transfer learning are usually used to counter the class imbalance challenge [154].
5.6 Explainability and Trustworthiness of DL Models
The explainability in DL, e.g., how GNNs and Transformers make a specific decision, is still an area of active research [155]. GNNExplainer [156], PGM-Explainer [157], and SubgraphX [158] are some attempts to explain the decision-making process of GNNs. The explainability methods for Transformers have been analyzed in [159]. Existing efforts and a roadmap to improve the trustworthiness of GNNs have been presented in the latest survey [160]. However, explainability and trustworthiness of multimodal GNNs and Transformers is an under-research challenge.
5.7 Oversmoothing in GNNs
One particular challenge in using GNNs is oversmoothing, which occurs when the GNN is trained for too many iterations, causing the node representations to become almost similar [58]. This leads to a loss of information, decrease in the model’s performance, and generalization [161]. Regularization techniques such as dropout, weight decay, skip-connection, and incorporation of higher-order structures, such as motifs and graphlets, have been proposed, but building deep architectures that can scale and adapt to varying structural patterns of graphs is still an open challenge.
@@ 5.8 Modality Collapse
Modality collapse is a phenomenon that occurs in MML, where a model trained on multiple modalities may become over-reliant on a single modality, to the point where it ignores or neglects the other modalities [162]. Recent work explored the reasons and theoretical understanding of modality collapse [163], but counter-actions need active investigation by the ML community.
@@ 5.9 Dynamic and Temporal Data
Dynamic and temporal data refers to the data that changes over time [58]. Tumor surveillance is a well-known technique to study longitudinal cancer growth over multiple data modalities [18]. Spatio-temporal methods such as multiple instance learning, GNNs, and hybrid of multiple models can capture complex change in the data relationships over time, but learning on multimodal dynamic data is an active area of research in oncology [164].
@@ 5.11 Other Challenges
MML requires extensive computational resources and time to train models on a variety of datasets and tasks. Robustness and failure detection [166] are critical aspects of MML, particularly in applications such as oncology. Uncertainty quantification techniques, such as Bayesian neural networks [167], are still under-explored avenues in the MML. By addressing these challenges, it is possible to develop MML models that are able to surpass performance offered by single-modality models.
@@ 6 Multimodal Oncology Data Sources
Building central archives to unify the different collections of oncology data requires a dedicated effort. We have compiled a non-exhaustive list of datasets from data portals shared by the National Institute of Health (NIH) and others. The purpose of this compilation is to provide ML oncology researchers with a unified source of data. This periodically updated collection is available at https://lab-rasool.github.io/pan-cancer-dataset-sources/.
@@ 7 Conclusion
Research in integrating data across a few modalities has already shown encouraging results. There is no unified framework available for scaling across all possible modalities of cancer. The convergence of individual methodologies and data across varying scales may hold vital clues in creating a unified view of the disease that is more prognostic, predictive, and insightful than the individual view or modality. In this survey, we reviewed the multimodal learning task on oncology data. The future resides in developing a deployment-ready, scalable deep learning framework with inherent uncertainty quantification, interpretability, and generalizability to integrate oncology data across multiple scales, modalities, and resolutions.

### v3

@@ 4 MML - Challenges and Opportunities
4.1 Large Amounts of High-quality Data
DL models are traditionally trained on large datasets with enough samples for training, validation, and testing, such as JFT-300M [161] and YFCC100M [162], which are not available in the cancer domain. For example, the largest genomics data repository, the Gene Expression Omnibus (GEO) database, has approximately 1.1 million samples with the keyword ‘cancer’ compared to 3 billion images in JFT-300M [163]. Annotating medical and oncology data is a time-consuming and manual process that requires significant expertise in many different areas of medical sciences. Factors like heterogeneity of the disease, noise in data recording, background, and training of medical professionals leading to inter- and intra-operator variability cause lack of reproducibility and inconsistent clinical outcomes [164].
@@ 4.2 Data Registration and Alignment
Data alignment and registration refer to the process of combining and aligning data from different modalities in a useful manner [165]. In multimodal oncology data, this process involves aligning data from multiple modalities such as CT, MRI, PET, and WSIs, as well as genomics, transcriptomics, and clinical records. Data registration involves aligning the data modalities to a common reference frame and may involve identifying common landmarks or fiducial markers. If the data is not registered or aligned correctly, it may be difficult to fuse the information from different modalities accurately [166].
@@ 4.3 Pan-Cancer Generalization and Transference
Transference in MML aims to transfer knowledge between modalities and their representations to improve the performance of a model trained on a primary modality [166]. Because of the unique characteristics of each cancer type and site, it is challenging to develop models that can generalize across different cancer sites. Furthermore, the models trained on a specific modality, such as radiology images, will not perform well with other imaging modalities, such as histopathology slides. Fine-tuning the model on a secondary modality, multimodal co-learning, and model induction are techniques to achieve transference and generalization [167]. To overcome this challenge, mechanisms for improved universality of ML models need to be devised.
@@ 4.4 Missing Data Samples and Modalities
The unavailability of one or more modalities or the absence of samples in a modality affects the model learning, as most of the existing DL models cannot process the “missing information”. This requirement, in turn, constrains the already insufficient size of datasets in oncology. Almost all publicly available oncology datasets have missing data for a large number of samples [163]. Various approaches for handling missing data samples and modalities are gradually gaining traction. However, this is still an open challenge [168].
@@ 4.5 Imbalanced Data
Class imbalance refers to the phenomenon when one class (e.g., cancer negative/positive) is represented significantly more in the data than another class. Class imbalance is common in oncology data [168]. DL models struggle to classify underrepresented classes accurately. Techniques such as data augmentation, ensemble, continual learning, and transfer learning are used to counter the class imbalance challenge [168].
4.6 Explainability and Trustworthiness
The explainability in DL, e.g., how GNNs and Transformers make a specific decision, is still an area of active research [169, 170]. GNNExplainer [171], PGM-Explainer [172], and SubgraphX [173] are some attempts to explain the decision-making process of GNNs. The explainability methods for Transformers have been analyzed in [174]. Existing efforts and a roadmap to improve the trustworthiness of GNNs have been presented in the latest survey [175]. However, the explainability and trustworthiness of multimodal GNNs and Transformers is an open challenge.
4.7 Over-smoothing in GNNs
One particular challenge in using GNNs is over-smoothing, which occurs when the GNN is trained for too long, causing the node representations to become almost similar [73]. This leads to a loss of information, a decrease in the model’s performance, and a lack of generalization [176]. Regularization techniques such as dropout, weight decay, skip-connection, and incorporating higher-order structures, such as motifs and graphlets, have been proposed. However, building deep architectures that can scale and adapt to varying structural patterns of graphs is still an open challenge.
@@ 4.8 Modality Collapse
Modality collapse is a phenomenon that occurs in MML, where a model trained on multiple modalities may become over-reliant on a single modality, to the point where it ignores or neglects the other modalities [177]. Recent work explored the reasons and theoretical understanding of modality collapse [178]. However, the counter-actions needed to balance model dependence on data modalities require active investigation by the ML community.
@@ 4.9 Dynamic and Temporal Data
Dynamic and temporal data refers to the data that changes over time [73]. Tumor surveillance is a well-known technique to study longitudinal cancer growth over multiple data modalities [12]. Spatio-temporal methods such as multiple instance learning, GNNs, and hybrid of multiple models can capture complex change in the data relationships over time; however, learning from multimodal dynamic data is very challenging and an active area of research [179].
@@ 4.11 Other Challenges
MML requires extensive computational resources to train models on a variety of datasets and tasks. Robustness and failure detection [181] are critical aspects of MML, particularly in applications such as oncology. Uncertainty quantification techniques, such as Bayesian neural networks [10], are still under-explored avenues in the MML. By addressing these challenges, it is possible to develop MML models that are able to surpass the performance offered by single-modality models.
@@ 5 Multimodal Oncology Data Sources
Building central archives to unify the different collections of oncology data requires a dedicated effort. We have compiled a non-exhaustive list of datasets from data portals shared by the National Institute of Health and others. The purpose of this compilation is to provide ML oncology researchers with a unified source of data. This periodically updated collection is available at https://lab-rasool.github.io/pan-cancer-dataset-sources/, [16].
@@ 6 Conclusion
Existing research into the integration of data across various modalities has already yielded promising outcomes, highlighting the potential for significant advancements in cancer research. However, the lack of a comprehensive framework capable of encompassing the full spectrum of cancer dataset modalities presents a notable challenge. The synergy between diverse methodologies and data across different scales could unlock deeper insights into cancer, potentially leading to more accurate prognostic and predictive models than what is possible through single data modalities alone. In our survey, we have explored the landscape of multimodal learning applied to oncology datasets and the specific tasks they can address. Looking ahead, the key to advancing this field lies in the development of robust, deployment-ready MML frameworks. These frameworks must not only scale efficiently across all modalities of cancer data but also incorporate capabilities for uncertainty quantification, interpretability, and generalizability. Such advancements will be critical for effectively integrating oncology data across multiple scales, modalities, and resolutions. The journey towards achieving these goals is complex, yet essential for the next leaps in cancer research. By focusing on these areas, future research has the potential to significantly enhance our understanding of cancer, leading to improved outcomes for patients through more informed and personalized treatment strategies.